For the hypothesis test H0: µ = 10 against H1: µ < 10 with variance unknown and n = 20, let the value of the test statistic be t0 = 1.25. Find the approximate the P-value.
Answer #1

solution'

given that

n =20

df =n - 1 = 20 - 1 = 19

t0 = 1.25

p value= 0.0086